Use slope-intercept form to write the equation of the line with the given properties.
m = 2, passing through (−2, −1)

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

point slope formula y-y1=m(x-x1)
y+1=2(x+2) watch the signs
y+1=2x+4
y=2x+3

Answer by ikleyn(52787)

.

At this site,  there is a group of lessons related to this class of problems
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point (*)
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
    - Advanced problems on finding equations for straight lines

    - OVERVIEW of lessons related to the slope of a straight line

The most relevant to your current problem is the lesson marked  (*)  in the list.
So start from this lesson.

But if you want to learn the subject in all its aspects,  you need to learn all these lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.